OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13)?

OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13)? 4x − y = 5 x − 4y = 5 x + 4y = 5 4x + y = 5